Circular No. 6561

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 BA_{6} (SPACEWATCH) MPEC 1997-CO5 and 1997-C13 give observations of what appeared

(using telescopes of aperture up to 1.2 m) to be an asteroidal object discovered by the Spacewatch program on Jan. 31 (with also prediscovery images on Jan. 11). The object, designated 1997 BA_{6}, was distant and intrinsically bright (r = 9.0 AU, H = 10.0).

J. W. Parker and H. F. Levison, Southwest Research Institute; and R. Fesen, Dartmouth College, report that 21 two-minute exposures with the M.I.T.-Dartmouth-Michigan 2.4-m Hiltner Telescope at Kitt Peak on Feb. 10 show 1997 BA_{6} clearly to have a coma. All the images were centroided, shifted and summed to produce an equivalent 42-min exposure wherein the object was stationary. The FWHM’s of comparison stars (similarly centroided, shifted and summed) were all consistently 1′′.2, while the FWHM of 1997 BA_{6} was 1′′.7.

K. Meech, O. Hainaut and J. Bauer, University of Hawaii, report: “CCD images of 1997 BA_{6}, obtained with the University of Hawaii 2.2m telescope on Mauna Kea on Feb. 17 and 18 UT, show a low-surface-brightness coma surrounding the object. The images were taken through the Kron-Cousins R filter, and the coma was visible in a 1200-s integration on the first night, while a deeper image (4800 s) was obtained on Feb. 18 in conditions of moderate seeing (0′′.9). The asymmetric coma extends to a distance of 3′′.7 (2.1 × 10^{4} km) in p.a. 7^{o}. The brightness falls off very rapidly from the core—by steeper than ρ^{−2}, ρ being the projected distance from the image center.”

The following orbital elements are taken from MPC 29067:

Epoch = 1999 Dec. 8.0 TT

T = 1999 Dec. 4.3655 TT  ω = 286.^{o}1624
e = 0.987450  Ω = 317.6522 }2000.0
q = 3.447727 AU  i = 72.4459
1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
Feb. 21 9^{h}13.^{m}28 +12^{o}58.^{′}5 7.929 8.885 164.^{o}1 1.^{o}8 19.5
Mar. 3 9 08.63 +12 58.8 7.926 8.821 153.0 2.9 19.5
13 9 04.40 +12 57.7 7.953 8.758 142.1 4.0 19.6
23 9 00.73 +12 54.8 8.004 8.695 131.3 4.9 19.6
Apr. 2 8 57.75 +12 49.6 8.077 8.631 120.8 5.7 19.7
12 8 55.54 +12 41.9 8.165 8.568 110.5 6.3 19.7
22 8 54.14 +12 31.5 8.264 8.504 100.4 6.7 19.7
1997 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden